Find the length of a diagonal of a square if its area is 162 square feet.

I don't even know how to set this problem up. Thanks for the help.

Answer by nerdybill(7384)

Find the length of a diagonal of a square if its area is 162 square feet.
.
Let x = side of the square
then
x^2 = 162
x = sqrt(162)
x = sqrt(2*9*9)
x = 9sqrt(2)
.
Let d = diagonal
then
(9sqrt(2))^2 + (9sqrt(2))^2 = d^2
2(9sqrt(2))^2 = d^2
2(81*2) = d^2
324 = d^2
sqrt(324) = d
18 feet = d